The Tennessee Jack Distillery produces custom-blended whiskey. A particular
blend consists of rye and bourbon whiskey. The company has received an order
for a minimum of 400 gallons of the custom blend. The customer specified
that the order must contain at least 40 percent rye and not more than 250
gallons of bourbon. The customer also specified that the blend should be
mixed in the ratio of two parts rye to one part bourbon. The distillery can
produce 500 gallons per week, regardless of the blend. The production
manager wants to complete the order in one week. The blend is sold for $12
per gallon. The distillery company’s cost per gallon is $4 for rye and $2
for bourbon. The company wants to determine the blend mix that will meet
customer requirements and maximize profits.

Answer by Edwin McCravy(20065)

>>>.....determine the blend mix that will meet
customer requirements and maximize profits.
Let x = the gallons of blend that will meet customer requirements.
The distillery can produce (no more than) 500 gallons per week, regardless
of the blend.
x ≤ 500

the blend should be mixed in the ratio of two parts rye to one part bourbon.
gallons of rye = (2/3)x
gallons of bourbon = (1/3)x
the order must contain at least 40 percent rye
This requirement is automatically fulfilled by the above because 2/3 is
about 66.7% which is more than 40%, and is certainly "at least 40%."
the order must contain.....not more than 250 gallons of bourbon.
(1/3)x ≤ 250
     x ≤ 750

This requirement is automatically fulfilled because x ≤ 500

The blend is sold for $12 per gallon. The distillery company’s cost per
gallon is $4 for rye and $2 for bourbon.
 

Profit = 12x - 4(2/3x) - 2(1/3x) = (26/3)x

Since (26/3)x is a increasing function, since its slope 26/3 is positive,
so the maximum profit will be greatest when x is greatest, which is when
x = 500, so we substitute:

Maximum profit = (26/3)(500) = $4333.33 when there is 500 gallons of custom-
blended whiskey.

Edwin